### v1

@@ 5.2 End-to-end Evaluation
Coding | 8A40 | TP=2, PP=1 | 4 Prefill Replicas |
@@
Conversation | 6A40 | TP=2, PP=1 | 3 Prefill Replicas |
@@ Appendix A NP-Hardness of Deployment Planning
Finding the optimal deployment plan that maximizes the overall SLO for deploying multiple models on a heterogeneous GPU cluster with variable interconnect topology and computational capabilities is non-trivial. In particular, we show that this problem is NP-hard by transforming it into the well-known NP-hard Job Shop Scheduling Problem (JSSP) Sotskov & Shakhlevich (1995), Omar et al. (2006).
@@ Appendix E Implementation Details
The workload profiler monitors the real-time workload patterns, including the average prompt length of incoming requests and average output length of generated responses. These patterns are utilized to analyze the prefill and decode cost for each single request. For instance, in contemporary LLM services, common workload scenarios include coding and conversation Patel et al. (2023), where both typically have a median prompt length exceeding 1000 tokens. However, the coding service produces much fewer output tokens, with a median of 13, while the conversation service generates a larger number of output tokens, with a median of 129. Undoubtedly, the overall system workload varies when the proportions of incoming requests for various services change in real-time. Once an obvious workload shift is detected, the workload profiler will notify the scheduler.
The task coordinator is in charge of the request dispatching among the prefill and decode replicas. Upon receiving a request, the task coordinator assigns the appropriate prefill replica and decode replica, respectively. The assignment is guided by the deployment plan generated by the scheduler. The task coordinator is mainly based on an open-source implementation of decentralized computation coordination Yao (2023) that utilizes libP2P LibP2P (2023) to establish connections among the work groups in a peer-to-peer network.

### v2

@@ 5.2 End-to-end Evaluation
Coding |
8A40 | TP=2, PP=1 | 4 Prefill Replicas |
@@
Conversation |
6A40 | TP=2, PP=1 | 3 Prefill Replicas |
@@ Appendix A NP-Hardness of Deployment Planning
Finding the optimal deployment plan that maximizes the overall SLO for deploying multiple models on a heterogeneous GPU cluster with variable interconnect topology and computational capabilities is non-trivial. In particular, we show that this problem is NP-hard by transforming it into the well-known NP-hard Job Shop Scheduling Problem (JSSP) sotskov1995np, omar2006job.
@@ Appendix E Implementation Details
The workload profiler monitors the real-time workload patterns, including the average prompt length of incoming requests and average output length of generated responses. These patterns are utilized to analyze the prefill and decode cost for each single request. For instance, in contemporary LLM services, common workload scenarios include coding and conversation patel2023splitwise, where both typically have a median prompt length exceeding 1000 tokens. However, the coding service produces much fewer output tokens, with a median of 13, while the conversation service generates a larger number of output tokens, with a median of 129. Undoubtedly, the overall system workload varies when the proportions of incoming requests for various services change in real-time. Once an obvious workload shift is detected, the workload profiler will notify the scheduler.
The task coordinator is in charge of the request dispatching among the prefill and decode replicas. Upon receiving a request, the task coordinator assigns the appropriate prefill replica and decode replica, respectively. The assignment is guided by the deployment plan generated by the scheduler. The task coordinator is mainly based on an open-source implementation of decentralized computation coordination yao2023open that utilizes libP2P libp2p to establish connections among the work groups in a peer-to-peer network.